OpenStudy (anonymous):

Choose the correct simplification of 3 over x to the power of negative 7. x to the 7th power over 3 3 over x to the 7th power 3x7 Already simplified.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\huge \frac{3}{x^{-7}}$$

OpenStudy (anonymous):

wht i dont get it?

OpenStudy (skullpatrol):

Is that the expression you want to simplify?

OpenStudy (anonymous):

I think so

OpenStudy (skullpatrol):

You should make sure that you have the right question first, no?

OpenStudy (anonymous):

can u just help me with the solution

OpenStudy (skullpatrol):

yes, if you make sure you have the right question please :-)

OpenStudy (anonymous):

yes i think i know the right question

OpenStudy (skullpatrol):

OK, let's start with what a negative exponent means.

OpenStudy (anonymous):

postive

OpenStudy (skullpatrol):

postive?

OpenStudy (anonymous):

7

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

so what does it means

OpenStudy (skullpatrol):

You said "7"

OpenStudy (skullpatrol):

I assume you know what you mean when you said it

OpenStudy (skullpatrol):

Maybe you just guessed at the answer?

OpenStudy (skullpatrol):

If you were guessing that the -7 becomes +7 then you are right.

OpenStudy (skullpatrol):

What did you base that guess on?

OpenStudy (anonymous):

no the positive 7 i didn't guess it

OpenStudy (skullpatrol):

$$\huge\frac {3}{x^{-7}}=?$$

OpenStudy (skullpatrol):

I'm only here to help you learn, I can't learn for you.

OpenStudy (anonymous):

so it \[\frac{ 3 }{ x ^{7} }\]

OpenStudy (skullpatrol):

Nope. Try to use what you know about negative exponents...

OpenStudy (skullpatrol):

$$\huge a^{-n}=\frac{1}{a^n}$$

OpenStudy (skullpatrol):

Have you seen this definition before?

OpenStudy (anonymous):

yes i think so

OpenStudy (skullpatrol):

Try to apply it to your question please.

OpenStudy (anonymous):

is it \[\frac{ x ^{7} }{ 3 }\]

OpenStudy (skullpatrol):

$$\huge\frac {3}{x^{-7}}=\frac{3}{\frac{1}{x^7}}$$

OpenStudy (skullpatrol):

Do you see how I used the definition on $$x^{-7}$$in the denominator?

OpenStudy (anonymous):

so its already simplified.

OpenStudy (skullpatrol):

Nope.

OpenStudy (anonymous):

wait its \[3x ^{7}\]

OpenStudy (skullpatrol):

Yes, by multiplying the numerator and the denominator by $$x^7$$ you get the simplification: $$3x^7$$

OpenStudy (anonymous):

so i was right

OpenStudy (anonymous):

so its c right just to make sure

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

thank u

OpenStudy (skullpatrol):

Thanks for trying to learn something new.

OpenStudy (anonymous):

okay